104.3 MHz (typical frequency for FM radio broadcasting) Express your answer using four significant figures.

Calculate the wavelength of each frequency of elec
Answer #1

MHZ= 106 Hz, 104.3 HZ= 104.3*106 HZ

from λν=c, c= 300000*1000 m/se=3*108 m/se

λ = c/ν=2.876 m

when ν=1080KHZ= 1080*103 HZ, λ=3*108/ (1080*1000)=277.7788m

when ν=835.6MHZ=835.6*106Hz, λ=3*108/(835.6*106)=0.359 m